# Synchronization of Dirac-Bianconi driven oscillators

### v1

@@ IV Phase description of Dirac-Bianconi driven oscillators
We now briefly discuss classical phase reduction theory, which is a method to reduce the dynamics of an oscillator to a single variable, the phase $\vartheta$, using a perturbative expansion valid under weak coupling Kuramoto_book ; nakao16 ; kuramoto2019concept ; monga2019phase . This method is widely applicable, since any self-sustained oscillator, i.e., with a stable limit cycle, can be reduced to a phase model if perturbations are small, and it has been recently extended to the case of (node-based) many-body interactions leon2025theory . In this section we will first introduce the phase reduction and then focus on the phase sensitivity function, which allows us to interpret the numerical results of the previous section.
@@ IV.1 Phase reduction approach
Because we are dealing with a population of oscillators, when uncoupled each subsystem displays an attractive limit cycle $\vec{X}_{c}(t)$. On the limit cycle we can define a phase variable $\vartheta$ with constant velocity $\dot{\vartheta}=\omega$. This definition can be extended to the whole basin of attraction with the concept of asymptotic phase and isochrons Win80 . This allows to assign a value of the phase to every point in the basin of attraction, $\vartheta_{j}(\vec{X}_{j})$. Applying the chain rule we confirm its velocity is constant:
@@
with $\vec{p}_{jk}(\vartheta_{j},\vartheta_{k})=\vec{g}_{jk}(\vec{X}_{j}(\vartheta_{j}),\vec{X}_{k}(\vartheta_{k}))$, and $Z_{j}(\vartheta_{j})=\nabla_{\vec{X}}\vartheta(\vec{X}_{c}(\vartheta))$ the phase sensitivity function, measuring phase response to infinitesimal perturbations winfree1967biological ; Win80 . Stated differently, the phase sensitivity function measures the response of the oscillator to the effect of the interactions. Note that we have neglected $\mathcal{O}(\epsilon^{2})$ terms  [^6] When close to the bifurcation point, which is not the case considered in this work, such terms may have some relevant effects leon19 ; leon22a ; leon20 ; bick2024higher ..
@@ IV.2 Phase sensitivity function
To numerically construct the phase model in Eq. (18), we need to determine the natural frequency $\omega$, the phase sensitivity function $Z(\vartheta)$, and the functional form of the limit cycle trajectory $\chi(\vartheta)$. These components are required to evaluate the interaction term $g_{jk}(\vec{X}_{j},\vec{X}_{k})$ and compute the coupling function $p_{jk}(\vartheta_{j},\vartheta_{k})$. However, analytical expressions for $Z(\vartheta)$ and $\chi(\vartheta)$ are typically available only for certain systems, such as the Stuart-Landau oscillator nakao16 , or the van der Pol oscillator Leon2023analytical .
In general, there are two main numerical approaches for computing the phase sensitivity function. The direct method, which involves applying small perturbations at different points along the limit cycle and measuring the resulting phase shifts. While conceptually straightforward, this method is computationally demanding, requiring careful control of the perturbation direction and magnitude to stay within the linear regime. A more elegant approach is given by the adjoint method, which consists in solving the adjoint equation ermentrout1996type ; hoppensteadt2012weakly ; brown2004phase , with the advantage of avoiding repeated perturbations. The adjoint equation is given by:
@@
ensuring that the phase grows uniformly with time along the limit cycle. The interested reader may consult nakao16 ; monga2019phase for a tutorial on this method and its implementation.
@@ IV.3 Phase sensitivity function of a Dirac-Bianconi driven oscillator
Figure 7: Phase sensitivity functions of the Dirac-Bianconi driven oscillator described in Eq. (9): we can see that the phase sensitivity functions of the variables on the nodes are small, while those of the variables on the links are larger. This explains why a Dirac-Bianconi coupling is necessary to achieve synchronization for weak coupling strengths. The parameters are the following: $a_{1}=a_{2}=0.7$, $a_{3}=-3$, $b_{1}=b_{2}=b_{3}=0.3$, $\delta_{1}=\delta_{2}=\delta_{3}=0.08$, $I_{1}=I_{2}=I_{3}=0.8$. Observe that the phase sensitivity function of $v_{3}$ is higher than the others, because $a_{3}$, which is the excitability in the FHN model, is higher. Due to the fact that $v_{3}$ interacts with $u_{1}$ and $u_{3}$, we see that also their phase sensitivity functions are higher than that of $u_{2}$, which is not directly interacting with $v_{3}$. With reference to Eq. (10) and to the Figs. of the previous section, note that the phase sensitivity functions of the two oscillators are similar, given that the two oscillators have similar parameters and frequencies.
@@
To understand the numerical results shown in the previous section, namely, why the two coupled oscillators synchronize better through a Dirac-Bianconi coupling, we compute the phase sensitivity functions of the two oscillators using the adjoint method described above. This method allows us to quantify how sensitive the phase of the oscillator is to small perturbations in each of its variables. The model setup and parameter values are the same as in the numerical example, so the two oscillators have similar dynamics and nearly identical frequencies. In Fig. 7, we display the phase sensitivity functions for all variables of one Dirac-Bianconi driven oscillator. It is clear that the phase sensitivity functions associated with the node variables are relatively small. On the other hand, the phase sensitivity functions associated with the link variables are larger. This indicates that interactions with respect to the link variables have a much stronger influence on the timing of the oscillations. This observation provides an explanation for why the Dirac-Bianconi coupling is better to achieve synchronization in this system, particularly when the coupling strength is weak. Since the link variables have a much greater effect on the phase, a coupling that acts through them is more effective at aligning the phases of the oscillators, thus yielding synchronization. A coupling acting only on the node variables is, indeed, insufficient, because their influence on the phase is too small. This is clearly shown in Fig. 5, where the coupling needs to be large in order for the frequencies to synchronize.
Finally, by computing the phase sensitivity functions for different Dirac-Bianconi systems, one can gain useful insights into which variables have a greater effect on the phase. This makes it possible to determine, in advance, the most suited coupling configuration to achieve synchronization. In this way, the phase sensitivity function serves as a practical tool for both guiding the design of effective coupling strategies and control the dynamics in systems of Dirac-Bianconi driven oscillators.
@@ V Conclusion
In this work we have introduced and described the dynamics of Bianconi-Dirac driven oscillators. These are higher-order networks in which the variables, called topological signals (or cochains), lie on the nodes, links, triangles, and higher-order simplicies and are coupled via the Dirac-Bianconi operator, which drives the system into an oscillatory behavior. We have fully described a simple Dirac-Bianconi driven oscillator, whose dynamics is inspired by the well-known FitzHugh-Nagumo neural model, in which the structure is solely made of nodes and links. We have put the fast variable, i.e., the voltage of the neuron, on the nodes, while the slow variable, i.e., the recovery variable (related to the ion density), on the links, and we have shown that, under certain values of the parameters, such structure exhibits a periodic dynamics. Then, by coupling two slightly different oscillators, we have shown that the classic diffusive coupling needs to be strong in order for the oscillators to synchronize, while, if the coupling is also mediated by the Dirac-Bianconi operator, synchronization is observed also for weak coupling. To understand this behavior, we derived a full phase description of a Dirac-Bianconi driven oscillator through phase reduction theory and obtained the phase sensitivity function, which measures the response of the phase to the interactions on nodes and links. We have shown that, for our example, the slow variables on the links have higher phase sensitivity functions than the fast variables on the nodes, meaning that a Dirac-Bianconi coupling is more fitting to achieve synchronization for weak values of the coupling. In future works, the concept of Dirac-Bianconi driven oscillator could be extended to directed simplicial complexes gong2024higher and higher-order multiplex networks krishnagopal2023topology , while different dynamics could be considered, such as stochastic systems zhu2022phase ; zhu2025complex or slow-fast systems whose phase sensitivity function could be approximated analytically Leon2023analytical . To achieve synchronization, optimization techniques could be applied with respect to the coupling connectivity nakao2021sparse , input function kato2021optimization , or coupling functions namura2024optimal , as well as control methods d2023controlling ; coraggio2025controlling . Our method, that we have made explicit for a simple case, can be straightforwardly extended to an arbitrary number of coupled Dirac-Bianconi driven oscillators of any dimensions and can provide a useful modeling tool in brain dynamics, where edge-signals are becoming increasingly relevant faskowitz2022edges .
@@ Acknowledgements
The work of R.M. is supported by a JSPS postdoctoral fellowship, grant 24KF0211. I.L. acknowledges support by Grant No. PID2021-125543NB-I00, funded by MICIU/AEI/10.13039/501100011033 and by ERDF/EU. Y.K. acknowledges JSPS KAKENHI JP22K14274 and JST PRESTO JPMJPR24K3 for financial support. H.N. acknowledges JSPS KAKENHI 25H01468, 25K03081, and 22H00516 for financial support. The authors are grateful to Ginestra Bianconi for discussions.

### v2

@@ IV Phase description of Dirac-Bianconi driven oscillators
We now briefly discuss classical phase reduction theory, which is a method to reduce the dynamics of an oscillator to a single variable, the phase $\vartheta$, using a perturbative expansion valid under weak coupling Kuramoto (1984); Nakao (2016); Kuramoto and Nakao (2019); Monga et al. (2019). This method is widely applicable, since any self-sustained oscillator, i.e., with a stable limit cycle, can be reduced to a phase model if perturbations are small, and it has been recently extended to the case of (node-based) many-body interactions León et al. (2025). In this section we will first introduce the phase reduction and then focus on the phase sensitivity function, which allows us to interpret the numerical results of the previous section.
@@ IV.1 Phase reduction approach
Because we are dealing with a population of oscillators, when uncoupled each subsystem displays an attractive limit cycle $\vec{X}_{c}(t)$. On the limit cycle we can define a phase variable $\vartheta$ with constant velocity $\dot{\vartheta}=\omega$. This definition can be extended to the whole basin of attraction with the concept of asymptotic phase and isochrons Winfree (1980). This allows us to assign a value of the phase to every point in the basin of attraction, $\vartheta(\vec{X}_{j})$. Applying the chain rule, we confirm its velocity is constant:
@@
with $\vec{p}_{jk}(\vartheta_{j},\vartheta_{k})=\vec{g}_{jk}(\vec{X}_{j}(\vartheta_{j}),\vec{X}_{k}(\vartheta_{k}))$, and $Z_{j}(\vartheta_{j})=\nabla_{\vec{X}}\vartheta(\vec{X}_{c}(\vartheta))$ the phase sensitivity function, measuring phase response to infinitesimal perturbations Winfree (1967, 1980). Stated differently, the phase sensitivity function measures the response of the oscillator to the effect of the interactions. Note that we have neglected $\mathcal{O}(\epsilon^{2})$ terms  [^7] When close to the bifurcation point, which is not the case considered in this work, such terms may have some relevant effects León and Pazó (2019, 2022, 2020); Bick et al. (2024)..
@@ IV.2 Phase sensitivity function
To numerically construct the phase model in Eq. (18), we need to determine the natural frequency $\omega$, the phase sensitivity function $Z(\vartheta)$, and the functional form of the limit cycle trajectory $\chi(\vartheta)$. These components are required to evaluate the interaction term $g_{jk}(\vec{X}_{j},\vec{X}_{k})$ and compute the coupling function $p_{jk}(\vartheta_{j},\vartheta_{k})$. However, analytical expressions for $Z(\vartheta)$ and $\chi(\vartheta)$ are typically available only for certain systems, such as the Stuart-Landau oscillator Nakao (2016), or the van der Pol oscillator León and Nakao (2023).
In general, there are two main numerical approaches for computing the phase sensitivity function. The direct method, which involves applying small perturbations at different points along the limit cycle and measuring the resulting phase shifts. While conceptually straightforward, this method is computationally demanding, requiring careful control of the perturbation direction and magnitude to stay within the linear regime. A more elegant approach is given by the adjoint method, which consists in solving the adjoint equation Ermentrout (1996); Hoppensteadt and Izhikevich (2012); Brown et al. (2004), with the advantage of avoiding repeated perturbations. The adjoint equation is given by:
@@
ensuring that the phase grows uniformly with time along the limit cycle. The interested reader may consult Nakao (2016); Monga et al. (2019) for a tutorial on this method and its implementation.
@@ IV.3 Phase sensitivity function of a Dirac-Bianconi driven oscillator
Figure 7: Phase sensitivity functions of the Dirac-Bianconi driven oscillator described in Eq. (9): we can see that the phase sensitivity functions of the variables on the nodes are small, while those of the variables on the links are larger. This explains why a Dirac-Bianconi coupling is necessary to achieve synchronization for weak coupling strengths. The parameters are the following: $a_{1}=a_{3}=0.7$, $a_{2}=-3$, $b_{1}=b_{2}=b_{3}=0.3$, $\delta_{1}=\delta_{2}=\delta_{3}=0.08$, $I_{1}=I_{2}=I_{3}=0.8$. Observe that the phase sensitivity function of $v_{3}$ is higher than the others, because $a_{3}$, which is the excitability in the FHN model, is higher. Due to the fact that $v_{3}$ interacts with $u_{1}$ and $u_{3}$, we see that also their phase sensitivity functions are higher than that of $u_{2}$, which is not directly interacting with $v_{3}$. With reference to Eq. (10) and to the Figs. of the previous section, note that the phase sensitivity functions of the two oscillators are similar, given that the two oscillators have similar parameters and frequencies.
@@
To understand the numerical results shown in the previous section, namely, why the two coupled oscillators synchronize better through a Dirac-Bianconi coupling, we compute the phase sensitivity functions of the two oscillators using the adjoint method described above. This method allows us to quantify how sensitive the phase of the oscillator is to small perturbations in each of its variables. The model setup and parameter values are the same as in the numerical example, so the two oscillators have similar dynamics and nearly identical frequencies. In Fig. 7, we display the phase sensitivity functions for all variables of one Dirac-Bianconi driven oscillator. It is clear that the phase sensitivity functions associated with the node variables are relatively small. On the other hand, the phase sensitivity functions associated with the link variables are larger. This indicates that interactions with respect to the link variables have a much stronger influence on the timing of the oscillations. This observation provides an explanation for why the Dirac-Bianconi coupling is better to achieve synchronization in this system, particularly when the coupling strength is weak. Since the link variables have a much greater effect on the phase, a coupling that acts through them is more effective at aligning the phases of the oscillators, thus yielding synchronization. A coupling acting only on the node variables is, indeed, insufficient, because their influence on the phase is too small. This is clearly shown in Fig. 5, where the coupling needs to be large in order for the frequencies to synchronize. Finally, by computing the phase sensitivity functions for different Dirac-Bianconi systems, one can gain useful insights into which variables have a greater effect on the phase. This makes it possible to determine, in advance, the most suited coupling configuration to achieve synchronization. In this way, the phase sensitivity function serves as a practical tool for both guiding the design of effective coupling strategies and control the dynamics in systems of Dirac-Bianconi driven oscillators.
IV.4 Phase coupling functions
Let us conclude by exploiting the phase reduction to further understand the synchronization dynamics of coupled Dirac-Bianconi driven oscillators and quantitatively explain the bifurcation causing the emergence of synchronization for strong enough couplings shown in Figs. 5 and 6.
After having performed the phase reduction, we obtain a phase equation for each of the two Dirac-Bianconi driven oscillators, DB1 and DB2, analogous to Eq. (19):
$\begin{cases}\dot{\vartheta}_{1}=\omega_{1}+\epsilon\vec{Z}_{1}(\vartheta_{1})\cdot\vec{p}_{12}(\vartheta_{1},\vartheta_{2}),\\ \dot{\vartheta}_{2}=\omega_{2}+\epsilon\vec{Z}_{2}(\vartheta_{2})\cdot\vec{p}_{21}(\vartheta_{2},\vartheta_{1}).\end{cases}$ (22)
As explained in Nakao (2016), under the hypothesis that $\omega_{1}-\omega_{2}\sim O(\epsilon)$, namely, $\omega_{1}=\omega+\Delta_{1}$ and $\omega_{2}=\omega+\Delta_{2}$, with $\Delta_{1,2}\sim O(\epsilon)$, we can introduce a slow phase for each DB oscillator, namely, $\varphi_{1}=\vartheta_{1}-\omega t$ and $\varphi_{2}=\vartheta_{2}-\omega t$. If we then describe the above system in terms of the slow phases, we obtain, after averaging approximation,
$\begin{cases}\dot{\varphi}_{1}=\Delta_{1}+\epsilon\Gamma_{12}(\varphi_{1}-\varphi_{2}),\\ \dot{\varphi}_{2}=\Delta_{2}+\epsilon\Gamma_{21}(\varphi_{2}-\varphi_{1}),\end{cases}$ (23)
where
$\Gamma_{ij}=\frac{1}{T_{i}}\int_{0}^{T_{i}}\vec{Z}_{i}(\varphi_{i}+\omega\tau)\cdot\vec{p}_{ij}(\varphi_{i}+\omega\tau,\varphi_{j}+\omega\tau)d\tau.$ (24)
Finally, defining $\phi=\varphi_{1}-\varphi_{2}$, and $\Delta_{1}-\Delta_{2}=\Delta\omega$, we obtain the equation of the phase difference
$\dot{\phi}=\Delta\omega+\epsilon\Big(\Gamma_{12}(\phi)-\Gamma_{21}(-\phi)\Big)=\Delta\omega+\epsilon\Gamma(\phi).$ (25)
Figure 8: Dynamics of the reduced phase models of two coupled Dirac-Bianconi driven oscillators: we show the dynamics of the relative phase of the two Dirac-Bianconi driven oscillators corresponding to the settings of Fig. 5 (left panel) and Fig. 6 (right panel). The magenta curves represent the frequency difference of the two uncoupled oscillators; the blue curves indicate the $\Gamma$ function for couplings such that synchronization is not achieved, i.e., no fixed point in the dynamics of the relative phases; the black curves represent the $\Gamma$ function for couplings such that synchronization is achieved, i.e., two fixed points, given by the intersection of the black and magenta curves, one stable (corresponding to the synchronous phase lock state, one unstable). We observe that, while the phase description accurately describes the case with Dirac-Bianconi coupling (right panel, with reference to Fig. 6), it only provides approximate results in the case with diffusing coupling (right panel, with reference to Fig. 5). Indeed, the phase reduction predicts synchronization for $\epsilon$ well below $0.2$, while direct numerical simulations of the system (Fig. 5) show that synchronization emerges above that coupling strength. The reason for this discrepancy is that the phase reduction is valid as long as the coupling is weak. This is why the description of the case with Dirac-Bianconi coupling is accurate (synchronization is obtained for $\varepsilon\gtrsim 0.006$), while the case with diffusive coupling not (synchronization is obtained for $\varepsilon\gtrsim 0.2$). The parameters are $\bar{a}_{1}=\bar{a}_{3}=\tilde{a}_{1}=\tilde{a}_{3}=0.7$, $\bar{a}_{2}=\tilde{a}_{2}=-3$, $\bar{b}_{1}=\bar{b}_{2}=\bar{b}_{3}=\tilde{b}_{1}=\tilde{b}_{2}=\tilde{b}_{3}=0.3$, $\bar{\delta}_{1}=\bar{\delta}_{2}=\bar{\delta}_{3}=\tilde{\delta}_{1}=\tilde{\delta}_{2}=\tilde{\delta}_{3}=0.08$, $\bar{I}_{1}=\bar{I}_{2}=\bar{I}_{3}=\tilde{I}_{1}=\tilde{I}_{2}=0.8$ and $\tilde{I}_{3}=0.4$.
The function $\Gamma$ allows us to describe the dynamics of the phase difference of the two oscillators, as shown in Fig. 8. In the figure, we depict the $\Gamma$ function for the case of diffusive coupling (left panel) and Dirac-Bianconi coupling (right panel). We observe that, when the coupling strength is high enough, $\Gamma$ intersects $\Delta\omega$ giving rise to two fixed points, see Eq.(25), one stable and one unstable. However, comparing Fig. 8 with Figs. 5 and 6, we see that, while the case of Dirac-Bianconi coupling of Fig. 6 is well described by the phase approximation of Fig. 8 (right panel), this approximation fails in describing the case where the coupling is diffusive (see Fig. 8, left panel and Fig. 5). In fact, from the phase description it would appear that a coupling strength of $\epsilon=0.15$ would be enough for the system to synchronize, while this is not the case. This discrepancy occurs because the coupling strength necessary to synchronize in the case with diffusive coupling is not small, hence, we are not in the range of validity of the phase reduction at first order. In fact, with higher values of the coupling strength, second-order effects are non-negligible and more accurate results can be obtained with a second-order phase reduction, as done in León and Pazó (2019); Erik T. K. Mau (2023); Bick et al. (2024). On the other hand, the Dirac-Bianconi coupling enables synchronization for much smaller values of the coupling parameters, allowing us to obtain an accurate description by means of the phase reduction.
@@ V Conclusion
In this work we have introduced and described the dynamics of Bianconi-Dirac driven oscillators. These are higher-order networks in which the variables, called topological signals (or cochains), lie on the nodes, links, triangles, and higher-order simplicies and are coupled via the Dirac-Bianconi operator, which drives the system into an oscillatory behavior. We have fully described a simple Dirac-Bianconi driven oscillator, whose dynamics is inspired by the well-known FitzHugh-Nagumo neural model, in which the structure is solely made of nodes and links. We have put the fast variable on the nodes, while the slow variable on the links, and we have shown that, thanks to the Dirac-Bianconi coupling, such structure can exhibit a periodic dynamics. Then, by coupling two slightly different oscillators, we have shown that the classic diffusive coupling needs to be strong in order for the oscillators to synchronize, while, if the coupling is also mediated by the Dirac-Bianconi operator, synchronization is observed also for weak coupling. To understand this behavior, we derived a full phase description of a Dirac-Bianconi driven oscillator through phase reduction theory and obtained the phase sensitivity function, which measures the response of the phase to the interactions on nodes and links. We have shown that, for our example, the slow variables on the links have higher phase sensitivity functions than the fast variables on the nodes, meaning that a Dirac-Bianconi coupling is more fitting to achieve synchronization for weak values of the coupling. In future works, more realistic models could be considered. For example, one could use this approach to model the dynamics of macro-region of the brains: each region could be represented by the nodes and their voltage by 0-cochains, while the links with their $1$-cochains would stand for the flow of current. One may also consider $2$-simplices (i.e., triangles) to model the dynamics of $2$-cochains, which would represent the magnetic field (a bivector Dorst et al. (2007)) induced by the flow of current. Moreover, the concept of Dirac-Bianconi driven oscillator could be extended to directed simplicial complexes Gong et al. (2024) and higher-order multiplex networks Krishnagopal and Bianconi (2023), while different dynamics could be considered, such as stochastic systems Zhu et al. (2022, 2025) or slow-fast systems whose phase sensitivity function could be approximated analytically León and Nakao (2023). To achieve synchronization, optimization techniques could be applied with respect to the coupling connectivity Nakao et al. (2021), input function Kato et al. (2021), or coupling functions Namura and Nakao (2024), as well as control methods D’Souza et al. (2023); Coraggio et al. (2026). Our method, that we have made explicit for a simple case, can be straightforwardly extended to an arbitrary number of coupled Dirac-Bianconi driven oscillators of any dimensions and can provide a useful modeling tool in brain dynamics, where edge-signals are becoming increasingly relevant Faskowitz et al. (2022).
@@ Acknowledgements
R.M. acknowledges JSPS KAKENHI 24KF0211 for financial support. I.L. acknowledges support by Grant No. PID2021-125543NB-I00, funded by MICIU/AEI/10.13039/501100011033 and by ERDF/EU. Y.K. acknowledges JSPS KAKENHI JP22K14274 and JST PRESTO JPMJPR24K3 for financial support. H.N. acknowledges JSPS KAKENHI 25H01468, 25K03081, and 22H00516 for financial support. R.M. is grateful to Future University Hakodate for the hospitality. The authors are grateful to Ginestra Bianconi for discussions and for the invitation to contribute to this special issue.
Authors contribution
Conceptualization: R.M., Y.K.; Methodology: R.M., I.L., H.N.; Formal Analysis and Investigation: R.M.; Visualization: R.M.; Numerical Implementation: R.M.; Writing—original draft preparation: R.M.; Writing—review and editing: all authors; Project Coordination: R.M.; Funding acquisition: Y.K., H.N.